OpenStudy (anonymous):

Someone throws a ball with initial speed v0 = 10 m/s at an angle θ = 30° above horizontal. It travels forward a exactly 10 m before hitting the ground. How long did it take to travel forward 10 meters? (Neglect air resistance.)

OpenStudy (anonymous):

Time period=\[2Vosin \Theta /g\]

OpenStudy (anonymous):

\[\sin \theta = t/Vi\]